OpenStudy (anonymous):

solve for x: x^2-2x-15=0 is ? please help

OpenStudy (zzr0ck3r):

two numbers that multiply to get -15 and add to get -2 are ?

OpenStudy (anonymous):

-5 and -3

OpenStudy (zzr0ck3r):

-5*-2=15 -5+-2=-7

OpenStudy (zzr0ck3r):

close...

OpenStudy (zzr0ck3r):

one of them should be negative. which one?

OpenStudy (anonymous):

-5 plus 2

OpenStudy (anonymous):

(x-5)(x+3)

OpenStudy (anonymous):

x=5 and x= -3

OpenStudy (anonymous):

solve step by step

OpenStudy (zzr0ck3r):

\(-5\) and \(3\) then we expand the middle term \(x^2-2x-15=0\) becomes \(x^2-5x+3x-15=0 \iff x(x-5)+3(x-5)=0\iff (x+3)(x-5)=0 \\ \iff x=-3 \ \text{or } x=5\)

OpenStudy (anonymous):

would you write it again